1. A lifeguard has been given 500m of rope to enclose a rectangular swim area designated for toddlers and children. The swimming area will be accessed from the beach; therefore only three sides must be roped in. Determine the dimensions of the rectangle that would maximize the swimming area.
1
Expert's answer
2021-11-10T14:36:14-0500

swimming area:

"S=xy"

perimeter:

"P=2(x+y)=500+x"


then:


"y=\frac{500+x}{2}-x"


"S=x(\frac{500+x}{2}-x)"


"S'(x)=\frac{500+x}{2}-x-x/2=250-x=0"


"x=250,y=125"


"S_{max}=250\cdot125=31250" m2